# Activating Distributed Visual Region within LLMs for Efficient and Effective Vision-Language Training and Inference

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) typically learn visual capacity through visual instruction tuning, involving updates to both a projector and their LLM backbones. Drawing inspiration from the concept of visual region in the human brain, we investigate the existence of an analogous visual region within LLMs that functions as a cognitive core, and explore the possibility of efficient training of LVLMs via selective layers tuning. We use Bunny-Llama-3-8B-V for detailed experiments and LLaVA-1.5-7B and LLaVA-1.5-13B for validation across a range of visual and textual tasks. Our findings reveal that selectively updating 25% of LLMs layers, when sparsely and uniformly distributed, can preserve nearly 99% of visual performance while maintaining or enhancing textual task results, and also effectively reducing training time. Based on this targeted training approach, we further propose a novel visual region-based pruning paradigm, removing non-critical layers outside the visual region, which can achieve minimal performance loss. This study offers an effective and efficient strategy for LVLM training and inference by activating a layer-wise visual region within LLMs, which is consistently effective across different models and parameter scales.
@@
Siyuan Wang [^1] footnotemark: 1, Dianyi Wang2thanks:  Equal contribution., Chengxing Zhou311footnotemark: 1, Zejun Li2, Zhihao Fan4, Xuanjing Huang2 Zhongyu Wei2 1University of Southern California, 2Fudan University, 3Sun Yat-sen University, 4Alibaba Inc. sw_641@usc.edu; dywang24@m.fudan.edu.cn; zhouchx33@mail2.sysu.edu.cn
@@ 1 Introduction
Large Vision-Language Models (LVLMs) Li et al. (2023); Zhu et al. (2023); Bai et al. (2023); Liu et al. (2024) have emerged as an increasing research interest for interpreting and interacting with the world through both visual and linguistic channels. Existing LVLMs generally utilize advanced Large Language Models (LLMs), like FlanT5 Chung et al. (2022) and Vicuna Chiang et al. (2023), as their cognitive core, and align visual features from visual encoders with LLMs’ knowledge and reasoning abilities. This alignment has demonstrated remarkable performance across diverse visual tasks Lu et al. (2022); Liu et al. (2023b); Fu et al. (2024).
@@
Figure 1: Left: Perplexity of LLaVA with selected layers (in parentheses) reverted to Vicuna parameters on visual and textual tasks. Arrows indicate perplexity increases relative to LLaVA (visual tasks) and Vicuna (textual tasks). Both results suggest layer redundancy in LVLMs and degraded linguistic capability. Right: Accuracy of LLaVA-1.5-7B when pruning certain layers based on angular distance scores Gromov et al. (2024).
@@
To this end, we embark on empirical experiments with Bunny-Llama-3-8B-V He et al. (2024) across diverse visual tasks. Our findings indicate that sparsely and uniformly distributed layers within LLMs are the optimal position for visual learning while simultaneously preserving textual performance. This strategic visual region selection also outperforms previous layer importance strategies. Notably, updating only 25% of layers achieves nearly 99% performance on visual tasks while effectively saving training time. We further validate this conclusion with LLaVA-1.5-7B and LLaVA-1.5-13B Liu et al. (2023a), demonstrating its consistent applicability across models and parameter scales. Specifically, we achieve time reductions of nearly 23% for LLaVA-1.5-7B and LLaVA-1.5-13B, and 12% for Bunny-Llama-3-8B-V.
Additionally, as shown in Figure 1 (right), we find that commonly used layer-pruning strategies are ineffective for LVLMs, with even minimal layer removal causing significant performance degradation. In response, we propose a visual region-based pruning paradigm that selectively prunes less-important layers outside the visual region after targeted training. Specifically, we follow the angular distance based layer importance strategy Gromov et al. (2024) outside the visual region, and experimental results demonstrate that our paradigm is effective to minimizes performance decline. Overall, our work highlights promising potential for more efficient LVLMs training and inference.
@@ 2.1 Model Architecture
Mainstream LVLMs consist of three components: a LLM, a visual encoder, and a projector or connection module, aim to effectively leverage the capabilities of both the pre-trained visual model and LLMs. The visual encoder extracts visual features from images, commonly utilizing pre-trained models such as CLIP ViT-L/14 Radford et al. (2021). The connection module then projects these extracted features into word embedding space understandable by LLMs, commonly employing techniques such as linear projection  Tsimpoukelli et al. (2021), Q-former Li et al. (2023), or cross-attention layers Alayrac et al. (2022). This enables LVLMs based on LLMs cores, like Vicuna Chiang et al. (2023), FlanT5 Chung et al. (2022), and LLaMA Touvron et al. (2023) to process visual information in a similar manner as text.
@@ 2.2 Model Training
The training process of LVLMs can be broadly divided into two phases: pre-training and supervised fine-tuning. Unlike LLMs, both the pre-training and supervised fine-tuning of LVLMs utilize supervised image-text pairs for visual instruction tuning. The pre-training stage primarily utilizes large-scale captioning instruction data, which requests the model to briefly describe images. This phase enables the model to interpret image content, usually with LLMs’ weights frozen and the visual encoder optionally updated. Some works such as Qwen-VL Bai et al. (2023), also expand the pre-training to incorporate additional tasks like visual question answering, and update the LLMs component accordingly. During the supervised fine-tuning stage, high-quality instruction data are employed to enhance the LVLMs’ ability in following diverse visual instructions and engaging in conversations. The visual encoder in this stage is typically kept static while the LLMs are tuned. During both stages, the projector is consistently updated, ensuring the model’s evolving ability to bridge visual and textual data.
@@ 3 Experimental Setup
In this study, we conduct empirical experiments on Bunny-Llama-3-8B-V to investigate our hypothesis regarding the existence of a specific visual region within LLMs (Sec. 4.1$\sim$ 4.3), and apply our findings on LLaVA-1.5-7B and LLaVA-1.5-13B to validate its general applicability across different model configurations (Sec. 5.1).
@@ 3.1 LVLM Implementation
We employ Bunny-Llama-3-8B-V for investigation experiments, which builds upon the 32-layer Llama3-8B Touvron et al. (2023), and LLaVA-1.5-7B/13B, built on the 32/40-layer Vicuna-1.5-7B/13B Chiang et al. (2023), for validation. Since the LLM components remain frozen during pre-training, our focus is on the supervised fine-tuning stage using 695K and 665K language-image instruction-following instances for Bunny and LLaVA, respectively. Considering computational constraints, we use the LoRA Hu et al. (2021) strategy, highlighting that our approach is complementary to other efficient training methods. Training was conducted with DeepSpeed Song et al. (2023) configured for zero3 optimization on 8×A800 GPUs. Additional implementation details are available in the Appendix.
@@ 3.2 Evaluation Tasks
Our investigation spans diverse visual perception and cognition tasks, to comprehensively evaluate models and thoroughly examine our hypothesis.
3.2.1 Visual perception tasks
Visual perception tasks assess models’ ability to interpret and understand surface-level visual features, such as object identification and scene recognition, mirroring human sensory perception process.
  • OCRVQA Mishra et al. (2019): VQA by reading text in images through optical character recognition (OCR). We adopt the accuracy calculation method from  Bai et al. (2023) on the test set, allowing a certain margin of error.
  • DocVQA Mathew et al. (2021): VQA by interpreting document images. For DocVQA’s validation set, we employ the same evaluation method and metric as in OCRVQA.
  • RefCOCOg Yu et al. (2016): A variant of RefCOCO Kazemzadeh et al. (2014) featuring more complex object referring expressions in COCO images Lin et al. (2014). We assess the reference expression generation task on the test set using Intersection over Union as the metric.
  • TDIUC Kafle and Kanan (2017): a VQA task suite across 12 categories, primarily perception tasks (e.g., object presence, counting, recognition) with some cognition tasks (e.g., positional reasoning, affordance, subordinate). Accuracy is measured on the validation set.
3.2.2 Visual cognition tasks
Contrary to perception tasks, cognition tasks require deeper reasoning based on visual stimuli, drawing on prior knowledge and advanced decision-making abilities learned within LLMs. This process resembles cognitive thinking and manipulation in human mental activities.
Model Version Perception Cognition Avg
OCRVQA DocVQA RefCOCOg TDIUC MMBench GQA ScienceQA TextVQA
All layers 64.26% 29.45% 50.12% 83.84% 74.74% 64.29% 79.28% 62.11% 63.51%
@@ 3.2.2 Visual cognition tasks
Sparse & Uniform 62.65% 29.51% 48.33% 83.68% 73.88% 63.68% 78.78% 62.43% 62.88%
Consecutive Lower 61.38% 22.47% 46.49% 83.27% 73.63% 62.33% 75.26% 62.26% 60.89%
Consecutive Lower-middle 62.54% 26.13% 48.17% 83.77% 72.51% 62.81% 77.14% 60.96% 61.75%
Consecutive Upper-middle 62.32% 28.06% 43.12% 83.40% 70.27% 61.28% 78.83% 59.33% 60.83%
Consecutive Top 60.48% 26.47% 39.92% 83.22% 67.96% 60.30% 77.54% 58.71% 58.08%
Hybrid Top-Lower 57.63% 29.76% 41.79% 83.26% 72.25% 62.71% 77.99% 62.74% 61.02%
@@
Image Attention Score 63.65% 24.53% 43.62% 83.90% 72.59% 62.82% 77.59% 61.99% 61.34%
Parameter Change Ratio 63.94% 26.94% 47.67% 83.88% 73.54% 63.21% 78.68% 61.73% 62.45%
Block Influence Score 62.38% 28.45% 46.37% 83.73% 71.13% 61.93% 77.34% 59.93% 61.41%
Multimodal BI Score 61.48% 28.80% 46.68% 83.74% 73.02% 63.23% 77.24% 62.23% 62.05%
Angular Distance 60.95% 27.71% 46.74% 83.49% 73.88% 62.11% 77.14% 62.76% 61.85%
@@
  • MMBench Liu et al. (2023b): a benchmark mainly for cognition tasks, with some fine-grained perception tasks requiring knowledge and reasoning. For simpler model variant comparison, we calculate accuracy on the dev subset instead of submitting to the evaluation server.
  • GQA Hudson and Manning (2019): a dataset featuring real-world visual reasoning and compositional question answering.
  • ScienceQA Lu et al. (2022): sourced from elementary and high school science curricula, requiring external knowledge and reasoning. We focus on questions with image context.
  • TextVQA Singh et al. (2019): a dataset requiring reasoning about text in images. For TextVQA, ScienceQA, and GQA, we use the LLaVA evaluation codes to measure accuracy.
@@ 4.1 Where are visual region layers located within LLMs for effective visual learning?
  • Heuristic Layer Selection (1) We intuitively hypothesize that tuning sparsely and uniformly distributed layers (0,4,8,12,18,22,26,30) preserves LLMs’ existing knowledge and reasoning abilities while enabling visual learning. (2) We experiment with tuning consecutive 8-layer blocks at different positions in LLMs: lower layers (0$\sim$7), lower-middle layers (8$\sim$15), upper-middle layers (16$\sim$23), and top layers (24$\sim$31), with the latter being a common practice of efficient downstream fine-tuning Liao et al. (2024). (3) We test a hybrid of lower and top layers (0$\sim$3, 28$\sim$31).
@@
Model Scale Perception Cognition Avg
OCRVQA DocVQA RefCOCOg TDIUC MMBench GQA ScienceQA TextVQA
32 layers 64.26% 29.45% 50.12% 83.84% 74.74% 64.29% 79.28% 62.11% 63.51%
16 layers 62.42% 26.43% 49.15% 84.04% 74.83% 64.10% 78.93% 62.96% 62.86% (98.98%)
8 layers 62.65% 29.51% 48.33% 83.68% 73.88% 63.68% 78.78% 62.43% 62.88% (99.00%)
6 layers 62.25% 29.76% 47.71% 84.01% 75.00% 62.93% 77.54% 62.92% 62.78% (98.85%)
4 layers 62.40% 28.89% 46.00% 83.99% 73.71% 62.66% 77.69% 62.74% 62.26% (98.03%)
2 layers 57.96% 28.49% 44.67% 83.15% 72.68% 61.00% 78.48% 60.35% 60.85% (95.81%)
1 layer 53.68% 24.33% 38.47% 82.92% 68.64% 59.19% 77.69% 58.32% 57.91% (91.18%)
@@
The results are shown in Table 1. We observe that tuning sparsely and uniformly distributed layers achieves the best overall performance across perception and cognition tasks, closely matching the all-layers upper bound. In contrast, consecutive layers generally underperform, likely due to limited diversity in similar representations across adjacent layers Kornblith et al. (2019), which hinders adaptability to various tasks. This further underscores the superiority of sparsely and uniformly distributed layers. Notably, tuning top layers yields the worst performance, deviating from the conventional practice in domain-specific fine-tuning, where the last few layers are typically adjusted for downstream tasks Liao et al. (2024). This highlights a significant distinction between adapting to new modalities and new downstream domains. Furthermore, while importance-based metrics are effective for layer pruning during LLMs inference, they are less effective than our empirically selected sparse and uniform layers for visual learning.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) typically learn visual capacity through visual instruction tuning, involving updates to both a projector and their LLM backbones. Inspired by the concept of a visual region in the human brain, we investigate the existence of an analogous visual region within LLMs that functions as a cognitive core, and explore the potential of efficient training of LVLMs via selective layers tuning. Using Bunny-Llama-3-8B-V for detailed analysis and other three LVLMs for validation across diverse visual and textual tasks, we find that selectively updating 25% of LLMs layers, when sparsely and uniformly distributed, can preserve nearly 99% of visual performance and maintain or improve textual task results, while effectively reducing training time. Based on this targeted training approach, we further propose a novel visual region-based pruning paradigm, removing non-critical layers outside the visual region, which can achieve minimal performance loss. This study offers an effective and efficient strategy for LVLM training and inference by activating a layer-wise visual region within LLMs, which proves consistently effective across different models.
@@
Siyuan Wang [^1] footnotemark: 1, Dianyi Wang2,3thanks:  Equal contribution., Chengxing Zhou411footnotemark: 1, Zejun Li2, Zhihao Fan5, Xuanjing Huang2 Zhongyu Wei2,3 1University of Southern California, 2Fudan University, 3Shanghai Innovation Institute, 4Sun Yat-sen University, 5Alibaba Inc. sw_641@usc.edu; dywang24@m.fudan.edu.cn; zhouchx33@mail2.sysu.edu.cn
@@ 1 Introduction
Large Vision-Language Models (LVLMs) Li et al. (2023c); Zhu et al. (2023); Bai et al. (2023); Liu et al. (2024) have emerged as an increasing research interest for interpreting and interacting with the world through both visual and linguistic channels. Existing LVLMs generally utilize advanced Large Language Models (LLMs), like FlanT5 Chung et al. (2022) and Vicuna Chiang et al. (2023), as their cognitive core, and align visual features from visual encoders with LLMs’ knowledge and reasoning abilities. This alignment has demonstrated remarkable performance across diverse visual tasks Lu et al. (2022); Liu et al. (2023b); Fu et al. (2024).
@@
Figure 1: Left: Perplexity of LLaVA with selected layers (in parentheses) reverted to Vicuna parameters on visual and textual tasks. Arrows indicate perplexity increases relative to LLaVA (visual tasks) and Vicuna (textual tasks). (1) Perplexity increases in textual tasks after multimodal training compared to the LLM backbone, indicating multimodal training compromises LLMs’ linguistic abilities. (2) Perplexity decreases in visual tasks reverting certain layers (e.g., reverting layers 16–23 or 24-31 in LLaVA), suggesting these layers are redundant. Right: Accuracy of LLaVA-1.5-7B when pruning certain layers based on angular distance scores Gromov et al. (2024).
@@
To this end, we embark on empirical experiments with Bunny-Llama-3-8B-V He et al. (2024) across diverse visual tasks. Our findings indicate that sparsely and uniformly distributed layers within LLMs are the optimal position for visual learning while simultaneously preserving textual performance. This strategic visual region selection also outperforms previous layer importance strategies. Notably, updating only 25% of layers achieves nearly 99% performance on visual tasks while effectively saving training time. We further validate this conclusion with LLaVA-1.5-7B, LLaVA-1.5-13B Liu et al. (2023a) and Bunny-Phi3-mini-4B-V, demonstrating its consistent applicability across varying models and parameter scales. Specifically, we achieve time reductions of nearly 23% for LLaVA-1.5-7B and LLaVA-1.5-13B, and 12% for Bunny-Llama-3-8B-V.
Additionally, as shown in Figure 1 (right), we find that commonly used layer-pruning strategies are ineffective for LVLMs, with even minimal layer removal causing significant performance degradation. In response, we propose a visual region-based pruning paradigm that selectively prunes less-important layers outside the visual region after targeted training. Specifically, we follow the angular distance based layer importance strategy Gromov et al. (2024) outside the visual region, and experimental results demonstrate that our paradigm is effective to minimizes performance decline. Overall, our work highlights promising potential for more efficient LVLMs training and inference. Notably, our approach is flexibly complementary to other efficient training techniques, such as LoRA, as demonstrated in our experiments.
@@ 2.1 Model Architecture
Mainstream LVLMs consist of three components: a LLM, a visual encoder, and a projector or connection module, aim to effectively leverage the capabilities of both the pre-trained visual model and LLMs. The visual encoder extracts visual features from images, commonly utilizing pre-trained models such as CLIP ViT-L/14 Radford et al. (2021). The connection module then projects these extracted features into word embedding space understandable by LLMs, commonly employing techniques such as linear projection  Tsimpoukelli et al. (2021), Q-former Li et al. (2023c), or cross-attention layers Alayrac et al. (2022). This enables LVLMs based on LLMs cores, like Vicuna Chiang et al. (2023), FlanT5 Chung et al. (2022), and LLaMA Touvron et al. (2023) to process visual information in a similar manner as text.
@@ 2.2 Model Training
The training of LVLMs can be broadly divided into two phases: pre-training and supervised fine-tuning. Unlike LLMs, both phases utilize supervised image-text pairs for visual instruction tuning. Pre-training primarily uses large-scale captioning instruction data, guiding the model to briefly describe images. This phase enables the model to interpret image content, usually with LLMs’ weights frozen and the visual encoder optionally updated. Some works such as Qwen-VL Bai et al. (2023), expand the pre-training to include additional tasks like visual question answering, updating the LLM component accordingly. Supervised fine-tuning employs high-quality instruction data to enhance the LVLMs’ ability to following diverse visual instructions and engaging in conversations. The visual encoder in this stage is typically kept static while the LLMs are tuned. During both stages, the projector is consistently updated, ensuring the model effectively bridges visual and textual data.
@@ 3 Experimental Setup
In this study, we conduct empirical experiments on Bunny-Llama-3-8B-V to investigate our hypothesis regarding the existence of a specific visual region within LLMs (Sec. 4.1$\sim$ 4.3), and apply our findings on LLaVA-1.5-7B, LLaVA-1.5-13B and Bunny-Phi3-mini-4B-V to validate its general applicability across different models (Sec. 5.1).
@@ 3.1 LVLM Implementation
We employ Bunny-Llama-3-8B-V for investigation, which builds upon the 32-layer Llama3-8B Touvron et al. (2023), and LLaVA-1.5-7B/13B, built on the 32/40-layer Vicuna-1.5-7B/13B Chiang et al. (2023), Bunny-Phi3-mini-4B-V based on 32-layer Phi-3-mini for validation. Since the LLM components remain frozen during pre-training, we focus on the supervised fine-tuning stage using 695K and 665K language-image instruction-following instances for Bunny and LLaVA. Considering computational constraints, we use LoRA Hu et al. (2021), highlighting that our approach is complementary to other efficient training methods. Additional implementation details are available in the Appendix.
@@ 3.2 Evaluation Tasks
Our investigation spans 10 visual tasks involving both perception and cognition, to comprehensively evaluate models and examine our hypothesis.
Visual perception tasks
assess models’ ability to interpret and understand surface-level visual features, like object identification and scene recognition, mirroring human sensory perception process. (1) OCRVQA Mishra et al. (2019): VQA by reading text in images through optical character recognition (OCR). We followBai et al. (2023) for accuracy calculation on the test set, allowing a margin of error. (2) DocVQA Mathew et al. (2021): VQA by interpreting document images. We use the same evaluation method and metric as OCRVQA on the validation set. (3) RefCOCOg Yu et al. (2016): A variant of RefCOCO Kazemzadeh et al. (2014) featuring more complex object referring expressions. We assess the reference expression generation on the test set using Intersection over Union metric. (4) TDIUC Kafle and Kanan (2017): covering 12 categories, primarily perception tasks (e.g., object presence, counting, recognition) with some cognition tasks (e.g., positional reasoning, affordance). Accuracy is measured on the validation set.
Visual cognition tasks
require deeper reasoning based on visual stimuli, drawing on prior knowledge and decision-making abilities learned within LLMs, mirroring human cognitive thinking and manipulation.
Model Version OCRVQA DocVQA RefCOCOg TDIUC MMBench GQA ScienceQA TextVQA MMMU SEED-IMG Avg
All layers 64.26% 29.45% 50.12% 83.84% 74.74% 64.29% 79.28% 62.11% 40.6% 73.13% 62.18%
@@
Sparse & Uniform 62.65% 29.51% 48.33% 83.68% 73.88% 63.68% 78.78% 62.43% 42.1% 72.61% 61.82%
Consecutive Lower 61.38% 22.47% 46.49% 83.27% 73.63% 62.33% 75.26% 62.26% 42.6% 72.66% 60.24%
Consecutive Lower-middle 62.54% 26.13% 48.17% 83.77% 72.51% 62.81% 77.14% 60.96% 38.8% 72.16% 60.50%
Consecutive Upper-middle 62.32% 28.06% 43.12% 83.40% 70.27% 61.28% 78.83% 59.33% 38.3% 70.45% 59.54%
Consecutive Top 60.48% 26.47% 39.92% 83.22% 67.96% 60.30% 77.54% 58.71% 37.0% 71.00% 57.26%
Hybrid Top-Lower 57.63% 29.76% 41.79% 83.26% 72.25% 62.71% 77.99% 62.74% 40.1% 72.59% 60.09%
@@
Image Attention Score 63.65% 24.53% 43.62% 83.90% 72.59% 62.82% 77.59% 61.99% 39.3% 72.29% 60.23%
Parameter Change Ratio 63.94% 26.94% 47.67% 83.88% 73.54% 63.21% 78.68% 61.73% 42.0% 72.85% 61.45%
Block Influence Score 62.38% 28.45% 46.37% 83.73% 71.13% 61.93% 77.34% 59.93% 38.9% 71.66% 60.18%
Multimodal BI Score 61.48% 28.80% 46.68% 83.74% 73.02% 63.23% 77.24% 62.23% 41.0% 72.25% 60.97%
Angular Distance 60.95% 27.71% 46.74% 83.49% 73.88% 62.11% 77.14% 62.76% 39.9% 73.01% 60.77%
@@
(5) MMBench Liu et al. (2023b): focuses on cognition tasks, with some fine-grained perception tasks requiring knowledge and reasoning. For model variant comparison, we report accuracy on the dev subset instead of submitting to the evaluation server. (6) GQA Hudson and Manning (2019): real-world visual reasoning and compositional question answering. (7) ScienceQA Lu et al. (2022): sourced from elementary and high school science curricula, requiring external knowledge and reasoning. We evaluate only image-based questions. (8) TextVQA Singh et al. (2019): requiring reasoning about text in images. (9) MMMU Yue et al. (2024): covering math, science, and commonsense reasoning with accuracy calculated. (10) SEED-IMG: The image-based QA from SEED-Bench Li et al. (2023a).
@@ 4.1 Where are visual region layers located within LLMs for effective visual learning?
  • Heuristic Layer Selection (1) We intuitively hypothesize that tuning sparsely and uniformly distributed layers (0,4,8,12,18,22,26,30) preserves LLMs’ existing knowledge and reasoning abilities while enabling visual learning. (2) We experiment with tuning consecutive 8-layer blocks at different positions in LLMs: lower layers (0$\sim$7), lower-middle layers (8$\sim$15), upper-middle layers (16$\sim$23), and top layers (24$\sim$31), with the latter being a common practice of efficient domain-specific fine-tuning Liao et al. (2024). (3) We test a hybrid of lower and top layers (0$\sim$3, 28$\sim$31).
@@
Model Scale OCRVQA DocVQA RefCOCOg TDIUC MMBench GQA ScienceQA TextVQA MMMU SEED-IMG Avg
32 layers 64.26% 29.45% 50.12% 83.84% 74.74% 64.29% 79.28% 62.11% 40.6% 73.13% 62.18%
16 layers 62.42% 26.43% 49.15% 84.04% 74.83% 64.10% 78.93% 62.96% 42.6% 72.75% 61.82%(99.42%)
8 layers 62.65% 29.51% 48.33% 83.68% 73.88% 63.68% 78.78% 62.43% 42.1% 72.61% 61.78%(99.36%)
6 layers 62.25% 29.76% 47.71% 84.01% 75.00% 62.93% 77.54% 62.92% 40.6% 72.67% 61.55%(98.99%)
4 layers 62.40% 28.89% 46.00% 83.99% 73.71% 62.66% 77.69% 62.74% 39.2% 72.14% 60.94%(98.01%)
2 layers 57.96% 28.49% 44.67% 83.15% 72.68% 61.00% 78.48% 60.35% 40.8% 72.35% 60.00%(96.49%)
1 layer 53.68% 24.33% 38.47% 82.92% 68.64% 59.19% 77.69% 58.32% 37.4% 70.69% 57.14%(91.89%)
@@
The results are shown in Table 1. We observe that tuning sparsely and uniformly distributed layers achieves the best overall performance across perception and cognition tasks, closely matching the all-layers upper bound. In contrast, consecutive layers generally underperform, likely due to limited diversity in similar representations across adjacent layers Kornblith et al. (2019), which hinders adaptability to various tasks. This further underscores the superiority of sparsely and uniformly distributed layers. Notably, tuning top layers yields the worst performance, deviating from the conventional practice in domain-specific fine-tuning, where the last few layers are typically adjusted for downstream tasks Liao et al. (2024). This highlights a significant distinction between adapting to new modalities and new downstream domains.
While some importance-based selections, such as parameter change ratio, yield close performance, all importance-based methods operate post-hoc that require a fully trained model to compute importance metrics for layer selection. This makes them primarily suitable for inference and applying them during LVLM training incurs significantly higher computational costs. In contrast, our heuristic method is training-free, allowing for greater flexibility and direct transferability across different models, enhancing its practical applicability. We compare importance-based selections to show that our sparsely and uniformly distributed layers even outperform these post-hoc strategies and also simplify the process.